OpenStudy (natsent06):

x/8 = 23/40, What is x?

OpenStudy (skullpatrol):

It is a letter

OpenStudy (anonymous):

x=23/5 multiply the x/8 by 5 to reach the LCM in the denominator of 40, and you see that 5x/40=23/40, or 5x=23, or x=23/5. You can either leave it as a fraction like that for a more exact answer or put it in a calculator and round your answer, depending on the instructions given in the question

OpenStudy (agent0smith):

To solve for x, multiply both sides by 8 Simplify.